OpenStudy (anonymous):

Alaine has 1 gallon of paint. She is going to pour it into a paint tray that measures 10 inches wide, 12 inches long, and 5 cm deep. (1 gallon = 231 in3, 1 inch = 2.54 cm) Which of the following scenarios will happen? The paint will not fill the tray by 369 cm3. The paint will not fill the tray by 5.22 in3. The paint will fill the tray exactly. The paint will overfill the tray by 5.22 in3.

OpenStudy (jdoe0001):

how many inches in 5cm?

OpenStudy (anonymous):

1.9

OpenStudy (jdoe0001):

so 1.9 inches so our tray in inches can hold up to 10 x 12 x 1.9 = \(\bf 228 in^3\) of paint so, now let's find out, how many cm in 10 inches, and 12 inches?

OpenStudy (anonymous):

in 10 there is 25 cm and in 12 there is 30

OpenStudy (jdoe0001):

ok, then our tray in in cm will hold up to 25 x 30 x 5 = \(\bf 3750 cm^3\) of paint so, Alice is about to dump in the tray 1gallon, that is \(\bf 231 in^3 \ or \ 587cm^3\) will it fit, overflow or not fit?

OpenStudy (jdoe0001):

hmmm, lemme correct something really quick

OpenStudy (anonymous):

ok

OpenStudy (jdoe0001):

your numbers are rounded up, so lemme not round them up and use the decimal values since I think they're expected

OpenStudy (anonymous):

okay so will it overflow

OpenStudy (jdoe0001):

\(\bf 10 \times 12 \times 1.968 = 236.22 in^3\\ 25.4 \times 30.48 \times 5 = 3870.96 cm^3\) Alice is about to pour \(\bf 231 in^3 \ or \ 586.74 cm^3\)

OpenStudy (anonymous):

586.74cm?

OpenStudy (jdoe0001):

ohhh smokes.... you're right.. the figure man! yes one sec

OpenStudy (anonymous):

Which of the following scenarios will happen? The paint will not fill the tray by 369 cm3. The paint will not fill the tray by 5.22 in3. The paint will fill the tray exactly. The paint will overfill the tray by 5.22 in3.

OpenStudy (jdoe0001):

\(\bf 231 \times 2.54^3 = 3785.41\\\quad \\ 10 \times 12 \times 1.968 = 236.22 in^3\\ 25.4 \times 30.48 \times 5 = 3870.96 cm^3\) she's about to pour \(\bf 231 in^3 \ or \ 3785.41 cm^3\)

OpenStudy (anonymous):

so which would it be

OpenStudy (jdoe0001):

well, obviously the tray can take more than 231 cubic inches, so it will fit notice that, the tray can hold 236.22 the paint is just 231 the difference is 5.22 the tray holds the paint, and still have more room for 5.22 cubic inches so \(\bf \text{The paint will not fill the tray by 5.22 in3.}\)

OpenStudy (anonymous):

okay thank you!

OpenStudy (jdoe0001):

yw